### v2

# arXiv:2602.21778v2[cs.CV] 27 Feb 2026

Abstract
Instruction-based image editing has achieved remarkable success in semantic alignment, yet state-of-the-art models frequently fail to render physically plausible results when editing involves complex causal dynamics, such as refraction or material deformation. We attribute this limitation to the dominant paradigm that treats editing as a discrete mapping between image pairs, which provides only boundary conditions and leaves transition dynamics underspecified. To address this, we reformulate physics-aware editing as predictive physical state transitions and introduce PhysicTran38K, a large-scale video-based dataset comprising 38K transition trajectories across five physical domains, constructed via a two-stage filtering and constraint-aware annotation pipeline. Building on this supervision, we propose PhysicEdit, an end-to-end framework equipped with a textual-visual dual-thinking mechanism. It combines a frozen Qwen2.5-VL for physically grounded reasoning with learnable transition queries that provide timestep-adaptive visual guidance to a diffusion backbone. Experiments show that PhysicEdit improves over Qwen-Image-Edit by 5.9% in physical realism and 10.1% in knowledge-grounded editing, setting a new state-of-the-art for open-source methods, while remaining competitive with leading proprietary models. All code, checkpoints, and datasets are available at https://liangbingzhao.github.io/statics2dynamics/.
Machine Learning, ICML
Figure 1: Bridging semantic alignment and physical plausibility. (Top) Despite high semantic fidelity, existing editing models frequently violate physical principles. (Bottom) Traditional image editing treats editing as a black box, learning a discrete mapping with underspecified constraints. Our approach reformulates editing as a Physical State Transition, leveraging continuous dynamics to constrain the state transition space from unreal hallucinations to physically valid trajectories.
1 Introduction
Instruction-based image editing, which aims to generate a new image following the given instruction, enables users to do visual creation much more easily. As user demands evolve from simple style transfer or object replacement to more complex scenarios involving hypothetical or counterfactual changes, the focus has shifted toward reasoning-based editing (huang2024smartedit; he2025reasoning; zhuo2025reflection). This demand has contributed to the emergence of unified multi-modal models (UMMs), such as Bagel (deng2025emerging) and Nano Banana (nanobanana), which leverage the inherent textual reasoning capabilities of Multimodal Large Language Model (MLLMs) to bridge the gap between abstract user intent and visual execution. While these unified frameworks have established a new backbone for image editing, their reasoning modules operate primarily at the semantic level rather than physical causality. Consequently, despite high semantic fidelity, state-of-the-art models frequently hallucinate artifacts that violate fundamental physical principles, as they prioritize object matching over physical plausibility.
This drawback becomes particularly apparent in scenarios governed by strict physical interactions. Consider a common scenario: inserting a straw into a transparent glass of water. As shown in Figure 1, while existing models can correctly identify the object (“straw”) and the location (“in the glass of water”), they frequently fail to render the optical refraction phenomenon, where the straw should appear disjointed or bent at the water surface. Instead, they tend to generate a naive straw’s position, maintaining geometric rigidity while violating optical physical laws. This discrepancy reveals a fundamental limitation: current models maximize semantic alignment at the cost of physical plausibility.
To bridge this gap, we argue that the next frontier in visual creation lies in physics-aware image editing: a paradigm where generated content must rigidly adhere to the causal rules of the physical world. To achieve this goal, we propose a fundamental shift in problem formulation: the editing process should not be viewed as a static mapping between independent images, but rather as a predictive physical state transition. Under this formulation, the source image represents an initial state, and the edit instruction specifies an external interaction or trigger that drives the scene toward a subsequent state under physical laws.
A central challenge is that standard paired images supervision in image editing provides only boundary conditions, leaving the transition itself underspecified. By contrast, temporal sequences show intermediate evidence of how states evolve, making video a natural source of supervision for learning transition priors. Therefore, we construct PhysicTran38K, a large-scale video-based dataset tailored to physical state transitions. Unlike existing benchmarks that predominantly emphasize semantic operations (e.g., add/remove/replace), PhysicTran38K is organized around interaction-driven triggers and law-governed transitions. We design hierarchical physics categories covering five major physical domains, 16 intermediate sub-domains, and spanning 46 distinct transition types. By employing a two-stage filtering pipeline, we finally get and annotate 38,000 high-quality transition data, providing explicit supervision for how physical states evolve over time.
However, leveraging video data introduces a practical mismatch: while videos supervise training, intermediate states are unavailable during inference. We address this train-test discrepancy by introducing PhysicEdit, a physics-aware editing framework built on Qwen-Image-Edit (wu2025qwen) that learns from video trajectories while remaining compatible with the single-image inference workflow. We propose a textual-visual dual-thinking mechanism that decouples physical understanding into two branches. First, a physically-grounded reasoning branch uses a frozen Qwen2.5-VL-7B (Qwen2.5-VL) to produce structured physical constraints as textual context. Second, an implicit visual thinking branch introduces learnable transition queries that are trained to learn transition priors from video. Concretely, intermediate keyframes provide supervision through two complementary encoders (DINOv2 (oquab2023dinov2) for structural semantics and a VAE (wu2025qwen) for fine-grained appearance), and the transition queries are aligned to these structure- and texture-level targets via dual projection heads. Finally, we align this guidance with diffusion’s coarse-to-fine generation through a timestep-aware modulation strategy that emphasizes structure at high noise and texture details at low noise.
In summary, our contributions are as follows:
  • We propose PhysicEdit, an end-to-end physics-aware editing framework with a textual-visual dual-thinking mechanism. It combines physically-grounded reasoning with implicit visual thinking to leverage transition priors learned from videos, enabling physically faithful edits.
  • We construct PhysicTran38K, a large-scale video-based dataset of approximately 38k video-instruction pairs organized by hierarchical physics categories.
  • Extensive experiments demonstrate that PhysicEdit achieves state-of-the-art performance among evaluated open-source models, while performing comparably to leading proprietary models, establishing a strong baseline for physics-aware image editing.
2 Related Works
Benefit from powerful diffusion models (ho2020denoising; songscore; bakrtoddlerdiffusion) in generating high-fidelity images, instruction-based image editing has made rapid progress. Early diffusion-based editors (hertz2022prompt; zhao2023cross) typically manipulate cross-attention or invert latents to trade off content preservation and edit strength, but often lack fine-grained controllability under complex structural changes. This motivates instruction-tuned approaches (brooks2023instructpix2pix; flux2024; wei2024omniedit; zhuo2025factuality; zhuo2025reflection) and, more recently, unified multimodal models for generalized instruction alignment. A representative example is Qwen-Image-Edit (wu2025qwen), which conditions an MMDiT (esser2024scaling) on frozen Qwen2.5-VL (Qwen2.5-VL) multimodal representations to replace standard text embeddings. In parallel, video data has been explored as a prior for improving editing. Earlier works (deng2025emerging; xiao2024omnigen; chen2025unireal) construct training pairs from video keyframes to enhance consistency, whereas recent efforts (wu2025chronoedit; rotstein2025pathways) shift toward generative reasoning. Notably, the concurrent work ChronoEdit (wu2025chronoedit) explicitly synthesizes intermediate frames as reasoning steps but incurs substantial computation and potential error accumulation. In contrast, we adopt an implicit paradigm that distills physical state transition priors into compact latent queries, enabling efficient feature-space dynamics simulation while inheriting physical fidelity from video data. A more detailed discussion of related work is provided in Appendix LABEL:sec:app_related.
Figure 2: Overview of the PhysicTran38K construction pipeline. Starting from hierarchical physics categories, we synthesize videos using Wan2.2-T2V-A14B, filtered by ViPE with an adaptive strategy to preserve high-dynamic transitions. Candidate videos conduct principle-driven verification by GPT-5-mini, adhering to a rigorous retention rule. Finally, Qwen2.5-VL-7B performs constraint-aware annotation, generating instructions and structured reasoning while incorporating verification results to prevent hallucinations.
3 Method
3.1 Problem Formulation
To bridge the gap between semantic alignment and physical fidelity, we first formalize the task of physics-aware image editing. Conventionally, instruction-based image editing is modeled as a discrete conditional mapping. Given a source image $I_{src}$ and an editing instruction $T_{edit}$, the model approximates a function $\mathcal{F}$:
$I_{tgt}=\mathcal{F}(I_{src},T_{edit}),$ (1)
where the goal is to generate a target image $I_{tgt}$ that follows $T_{edit}$ while preserving relevant content from $I_{src}$. While effective for semantic edits (e.g., changing a dog to a cat), this formulation treats the transformation as a black-box pixel update and does not explicitly model the underlying dynamics that govern how the scene should evolve.
[…]
$S_{final}=S_{0}+\int_{0}^{\tau}\Phi(S_{t},T_{edit};\Omega)\,dt,$ (2)
where $\Phi$ denotes the state transition dynamics and $\tau$ is the duration of the interaction. The desired target image $I_{tgt}$ is the visual outcome of the accumulated state $S_{final}$.
The fundamental challenge arises because standard image editing dataset provide only the boundary conditions, i.e., $(I_{src},I_{tgt})$, leaving the transition dynamics $\Phi$ completely underspecified. As illustrated in Figure 1, for a given pair $(I_{src},T_{edit})$, there may exist multiple visually plausible endpoints, yet only a subset results from a valid physical trajectory under $\Omega$. Without constraints on the intermediate integral, models tend to violate specific physical laws that satisfy the instruction. We therefore leverage videos as supervision, since they expose intermediate states and directly constrain how states evolve over time, motivating our construction of PhysicTran38K (Section 3.2). During inference, however, intermediate states are unavailable, so we propose PhysicEdit (Section 3.3) to distill these transition priors from videos into a latent representation, effectively guiding the generation along a physically valid trajectory.
Figure 3: Overview of the PhysicEdit framework. (a) Training: We distill physical transition priors from video data into learnable transition queries. These queries are supervised by complementary visual features extracted from intermediate keyframes. (b) Inference: PhysicEdit follows a sequential workflow. The frozen MLLM first generates physically-grounded reasoning, which is then concatenated with the learned transition queries to serve as the condition for the diffusion backbone.
3.2 Physics-Driven Data Construction
We construct PhysicTran38K, a video-based dataset for physics-aware image editing, by casting editing as physical state transitions. As illustrated in Figure 2, our data construction pipeline is designed to transform hierarchical physics categories into high-quality video data through three stages: structured generation with a video generation model, camera movement and principle-driven verification, and constraint-aware reasoning generation. We provide all the system prompts in Appendix LABEL:sec:sys_prompts.
Hierarchical Physics Categories.
We begin by establishing a comprehensive physics taxonomy to ensure broad coverage of physical dynamics. As shown in Figure 2(a), we organize physical laws into five primary state domains: Mechanical, Thermal, Material, Optical, and Biological. Under these domains, we define 16 intermediate sub-domains and 46 transition types (e.g., refraction, melting, germination), which serve as the specific physical laws $\Omega$ governing our dataset. For each transition, we curate specialized object pools containing entities that naturally exhibit the corresponding physical changes.
Structured Generation.
Leveraging the transition categories, we synthesize videos utilizing a structured generation pipeline. We utilize GPT-5-mini to sample objects from the object pool and to instantiate a fixed Wan Prompt template used for video generation: [Start State] + [Trigger Event] + [Transition Description] + [Final State]. Using Wan2.2-T2V-A14B (wan2025wanopenadvancedlargescale), we generate 1,0001{,}000 raw videos per transition type. We include a static-camera constraint in the prompt so that pixel variations primarily reflect state $\{S_{t}\}$ transition rather than viewpoint changes. As shown in Figure 2(b), this pipeline successfully synthesizes diverse high-fidelity transitions across all defined physical domains.
Camera Movement and Principle-driven Verification.
As illustrated in Figure 2(c), raw generations undergo a two-stage filtering process to ensure both viewpoint stability and physical correctness. First, we notice that current text-to-video models’ instruction following ability is not perfect, leading to unwanted viewpoint shifts. We therefore apply ViPE (huang2025vipe) as a geometric stability filter. In practice, ViPE can confuse large non-rigid deformations with viewpoint changes. To reduce false rejections, we adopt an adaptive strategy that relaxes the ViPE threshold for transition types expected to induce substantial deformation.
Next, we verify whether the video is consistent with its intended physical laws. Given the Wan Prompt and transition type, we prompt GPT-5-mini to propose $N$ (typically $3$) transition-specific principles and classify them as align, contradict, or unknown based on the keyframes. We then calculate a verification score $S_{\text{verify}}$ and adopt a rigorous retention rule, preserving the video only if $S_{\text{verify}}=N_{\text{align}}/N_{\text{total}}\geq 0.5$. Crucially, rather than simply discarding flawed parts, we record contradicted principles as negative evidence to be used as constraints during the reasoning generation phase.
Constraint-Aware Reasoning Generation.
Finally, we convert each retained video into training supervision that matches our formulation in Section 3.1. From each clip, we extract the first and last frames to form an editing pair $(I_{src},I_{tgt})$ We also sample intermediate keyframes along the video trajectory to serve as the ground truth for our visual guidance mechanism in Section 3.3. We then use Qwen2.5VL-7B (Qwen2.5-VL) to generate an editing instruction $T_{edit}$ describing the trigger event and brief desired evolution, along with a structured transition reasoning that summarizes observable evidence of the initial state $S_{0}$, middle transition process $S_{1}...S_{T-1}$, and the final state $S_{T}$. To ensure text-visual consistency, we integrate the verification outcomes as logical constraints during annotation: principles classified as align are encouraged as mechanisms consistent with $\Omega$, while contradicted or unknown principles are treated as hard negative constraints and must be explicitly excluded. This constraint-aware annotation mitigates over-optimistic physical descriptions when samples contain artifacts, allowing us to retain more usable videos and yielding approximately 38,00038{,}000 samples.
3.3 PhysicEdit Framework
We present PhysicEdit, an end-to-end physics-aware editing framework built upon the Qwen-Image-Edit backbone. To bridge the gap between static image editing and dynamic physical laws, we instantiate a textual-visual dual-thinking mechanism with two complementary branches: physically-grounded reasoning, which uses a frozen Qwen2.5-VL-7B to provide explicit physical constraints as textual context, and implicit visual thinking, which introduces learnable transition queries to implicitly reconstruct transition priors in the latent space.
Physically-Grounded Reasoning.
This branch produces a detailed physics reasoning that equips the following editing with explicit constraints before visual generation. Given the source image and the editing instruction, the frozen Qwen2.5-VL generates a structured physics reasoning trace that describes: (i) what physical laws and constraints must be respected, (ii) how the change should unfold causally in the scene, and (iii) how the relevant materials should behave throughout the transition. During training, we use the precomputed constraint-aware reasoning annotations from PhysicTran38K and ask Qwen2.5-VL to generate the reasoning trace on the fly during inference. This step serves as contextual grounding for the subsequent generation, while requiring no parameter updates to the MLLM.
Implicit Visual Thinking.
While text reasoning provides logical constraints, precise physical rendering requires visual transition priors that are not explicitly observable from the input image or text reasoning. To avoid modeling entire video frames which is cumbersome, we introduce Implicit Visual Thinking, as illustrated in Figure 3.
Inspired by MetaQuery (pan2025transfer), we append $K$ (typically 64) learnable transition queries immediately after the reasoning prefix. These queries are processed by the frozen Qwen2.5-VL together with the text and source image, producing context-aware query embeddings. During training, we supervise the query embeddings with pseudo-target transition features extracted from the associated video in PhysicTran38K. Specifically, we sample intermediate keyframes and encode them using two frozen visual encoders: DINOv2 for semantic structure and the Qwen-Image-Edit VAE for fine-grained texture. We then compress the encoder outputs into fixed-length pseudo-target embeddings, denoted as $F_{\text{DINO}}$ and $F_{\text{VAE}}$, using learnable feature extractors. In parallel, we map the transition-query embeddings through two projection heads to obtain the corresponding predictions, $\hat{F}_{\text{DINO}}$ and $\hat{F}_{\text{VAE}}$. We train the transition queries by aligning these predictions to the pseudo-targets. In this way, the transition queries learn to implicitly represent the missing evolution between $I_{src}$ and $I_{tgt}$, so that at inference time they can be instantiated solely from the source image, the instruction, and the physically-grounded reasoning. For more details, please refer to Appendix LABEL:sec:imple_details.
Timestep-Aware Dynamic Modulation.
Since diffusion models typically follow a coarse-to-fine trajectory, where global structure is formed before local texture is refined, we mix structure- and texture-level transition guidance according to the diffusion timestep. Let $t\in[0,1]$ denote the diffusion timestep, where larger $t$ corresponds to higher noise. During training, we construct the transition guidance from the video-derived pseudo-targets:
$\small F_{\text{tran}}^{(t)}=t\cdot F_{\text{DINO}}+(1-t)\cdot F_{\text{VAE}}.$ (3)
When $t\rightarrow 1$, the guidance emphasizes $F_{\text{DINO}}$ to provide structurally consistent cues; when $t\rightarrow 0$, it emphasizes $F_{\text{VAE}}$ to refine texture-level details. $F_{\text{tran}}^{(t)}$ is then sent to MMDiT as condition. At inference time, the video-derived pseudo-targets are unavailable, and we instead instantiate the guidance from the transition queries’ predictions:
$\small\hat{F}_{\text{tran}}^{(t)}=t\cdot\hat{F}_{\text{DINO}}+(1-t)\cdot\hat{F}_{\text{VAE}}.$ (4)
The timestep-conditioned transition feature is injected into the diffusion model together with the physically-grounded reasoning condition to guide physics-aware generation.
Table 1: Quantitative comparison on PICABench-Superficial evaluated by GPT-5 for instruction-based editing models, where LP, LSE, RFL, RFR, DFM, CSL, GST, LST denote Light propagation, Light Source Effects, Reflection, Refraction, Deformation, Causality, Global State Transition, Local State Transition, respectively.